OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes though the point (0, 2). Answer y = 5x + 2 y = -2/5x + 2 y = 2x + 5 y = 1/2x + 5

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

That is wrong, it is definitely not the third option.

OpenStudy (anonymous):

Are you familiar with the general form of the slope-intercept equation for a line and what its variables mean?

OpenStudy (anonymous):

No!

OpenStudy (anonymous):

Then that is a good place to start. The general form is y= mx + b "m" is the slope of the line and "b" is the y-intercept which is where the line crosses the y-axis. It has coordinates (0, b). It is necessary to know some basics so that you can learn from this problem how to do other similar problems on your own. To get the slope of a perpendicular line, you have to get the negative reciprocal of the first line. That means that if the first line has slope a/b the second line will have slope -b/a because (a/b)(-b/a) = -1 and this what you are in part looking for.

OpenStudy (anonymous):

Whats the anwser to my Q?

OpenStudy (anonymous):

The answer to your question is going to be arrived at mainly by you with me only to help you along. That is what this whole website is about. Learning. You learn and do the work. I help. That's how it works. The first line has slope 5/2, what is the slope of the second line which is perpendicular?

OpenStudy (anonymous):

Your first task is to find the second slope, "m", where m times 5/2 = -1 Use what I said about the negative reciprocal. Quid pro quo. That means that: Now it's your turn